OpenStudy (hellomynameiskam):

-24 + 33 + (-6) =???? show how u got it

OpenStudy (jackthegreatest):

33-24= 9

OpenStudy (jackthegreatest):

9-6= 3

OpenStudy (jackthegreatest):

Ur answer is 3

OpenStudy (hellomynameiskam):

another one : (1 + (-15) + -35 = ??

OpenStudy (hellomynameiskam):

@jackthegreatest